Eli's sister Emma is 6 years older than he is. His father is 3 times as old as Emma. The sum of their ages is 69. How old is each?
Answer by algebra hello(55)

Let eli's age=x
emma's age=x 6
father's age= 3(x 6)
according to question
x x 6 3(x 6)=69
5x 24=69
5x=69-24
x=45/5
x=9
Eli's age=9
Emma's age=9 6=15
Father's age=3*15=45